# Geometric Scaling of Bayesian Inference in LLMs

### v2

@@ Abstract.
Recent work has shown that small transformers trained in controlled “wind-tunnel” settings can implement exact Bayesian inference, and that their training dynamics produce a geometric substrate—low-dimensional value manifolds and progressively orthogonal keys—that encodes posterior structure. We investigate whether this geometric signature persists in production-grade language models. Across Pythia, Phi-2, Llama-3, and Mistral families, we find that last-layer value representations organize along a single dominant axis whose position strongly correlates with predictive entropy, and that domain-restricted prompts collapse this structure into the same low-dimensional manifolds observed in synthetic settings.
To probe the role of this geometry, we perform targeted interventions on the entropy-aligned axis of Pythia-410M during in-context learning. Removing or perturbing this axis selectively disrupts the local uncertainty geometry, whereas matched random-axis interventions leave it intact. However, these single-layer manipulations do not produce proportionally specific degradation in Bayesian-like behavior, indicating that the geometry is a privileged readout of uncertainty rather than a singular computational bottleneck. Taken together, our results show that modern language models preserve the geometric substrate that enables Bayesian inference in wind tunnels, and organize their approximate Bayesian updates along this substrate.
@@ 1. Introduction
Recent work (Paper 1 of this trilogy (Aggarwal et al., 2025a)) demonstrated that small transformers trained in controlled “Bayesian wind tunnels” perform exact Bayesian inference on bijection-learning and HMM-filtering tasks: posterior entropy, KL divergence, and predictive distributions match analytic solutions to within 0.1 bits. Paper 2 showed that these behaviors arise from specific geometric mechanisms created by gradient dynamics: value manifolds ordered by entropy, orthogonal key frames defining hypothesis directions, and layerwise attention sharpening implementing a geometric Bayes rule.
These findings established that transformers can implement Bayesian inference when trained on tasks with known posteriors. What remains open is whether the same geometric mechanisms persist in large, naturally trained LLMs, where ground-truth posteriors are unavailable.
@@
Throughout this trilogy, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—e.g., filtering posteriors over hidden states—not a posterior over network weights. This is a statement about the function the transformer computes, not about weight-space uncertainty.
Several factors make this question non-trivial:
@@
  • production models employ architectural optimizations (GQA, RoPE, sliding-window attention, MoE routing) absent from wind-tunnel setups;
@@
Rather than attempting to define Bayesian posteriors for natural language, we test whether the geometric substrate identified in Papers 1—-2 persists across architectures and training regimes. We treat these geometric signatures as invariants: if transformers rely on similar computational principles at scale, the same value—-key—-attention geometry should appear even when exact posteriors cannot be measured.
@@
First, domain restriction produces a decisive bridge. Under mixed-domain prompts, value manifold dimensionality varies substantially across architectures (PC1+PC2 ranging from $\sim$15% in Mistral to $\sim$99% in Pythia-410M), reflecting different inductive biases and training regimes. However, single-domain prompts consistently collapse the manifold toward one dimension (PC1+PC2 $\approx$ 70–95%), approaching the geometric regime observed in wind-tunnel experiments. This collapse shows that production LLMs contain the same entropy-ordered Bayesian axis that wind-tunnel transformers learn explicitly.
Second, Bayesian updating persists at inference time. In a controlled in-context learning experiment (SULA), models move smoothly along their value manifold as more evidence is supplied, and manifold coordinates correlate strongly with analytical Bayesian entropy. This demonstrates that the geometry is not merely a training artifact - it is used during inference.
Third, static and dynamic geometric signatures separate cleanly. Value manifolds and key orthogonality are universal across architectures, including sliding-window and MoE variants. Dynamic attention focusing, however, depends on routing capacity: strong in full-sequence MHA, moderate in GQA, and weak or noisy in Mistral. This matches the frame—-precision dissociation predicted in Paper 2.
@@
Papers 1 and 2 in this series established that transformers can implement exact Bayesian inference in controlled “wind—-tunnel” settings, and that gradient dynamics generically sculpt the value and key spaces into geometric substrates that support such inference. However, these results were obtained in synthetic domains with analytically specified likelihoods. The present work answers a distinct question: do production-grade LLMs, trained on heterogeneous natural language, spontaneously develop the same geometric signatures, and do these signatures track evidence integration during inference? We show that the three hallmarks of Bayesian geometry—-low-dimensional value manifolds, orthogonal hypothesis frames, and evidence-dependent movement along entropy-aligned directions—-persist across four model families (Pythia, Phi-2, Llama-3, Mistral). Moreover, we provide the first large-scale evidence that these structures are functionally engaged during inference in a naturalistic task (SULA), even when their causal role is distributed rather than bottlenecked. This establishes that Bayesian geometry is a stable inductive bias of modern transformers, not an artifact of synthetic tasks.
@@
    Domain-restriction bridge. When prompts are restricted to a coherent domain, the value manifold collapses to one or two principal components explaining 80–95% of variance, numerically matching the geometric regime predicted by Paper 1 and derived in Paper 2.
@@ 2. Background: Bayesian Geometry in Controlled Settings
Papers 1 and 2 established that small transformers trained in controlled “Bayesian wind-tunnel” settings perform near-exact Bayesian inference on tasks with analytically tractable posteriors. We summarize only the key findings needed for the production-model analysis; full experimental details appear in those papers.
@@
Two families of synthetic tasks provided ground-truth posteriors:
(1) Bijection learning (Paper 1): models infer a random bijection $\pi:V\to W$ from in-context examples. Because the hypothesis space has size $K!$ (e.g., $3.6\times 10^{6}$ for $K=10$), memorization is impossible; exact analytic posterior trajectories are available. Transformers achieve $\text{MAE}<0.1$ bits between model and Bayes-optimal predictive entropy.
(2) HMM filtering (Paper 1): models track posterior distributions over latent states in Hidden Markov Models with $S=5$ states and $V=5$ emissions. Transformers match analytic posteriors with KL divergence <0.05<0.05 bits, including strong length generalization.
@@
Paper 2 showed that cross-entropy gradients generate this geometry via coupled specialization of queries, keys, and values. A predicted frame—-precision dissociation emerges: attention patterns (the frame) stabilize early, while value manifolds (precision) continue refining.
@@
In this paper, we evaluate whether the static and dynamic geometric signatures identified in wind-tunnel models - entropy-ordered value manifolds, orthogonal hypothesis frames, and layerwise attention refinement - persist in production-scale LLMs and whether these structures are used during inference. Section Section 5 presents our empirical findings across architectures.
@@ 4.1. Model Selection
We selected three production models to span architectural and training variations:
Pythia-410M (Biderman et al., 2023): GPT-NeoX architecture, 24 layers, 16 attention heads, 1024 hidden dimensions. Trained on the Pile corpus (800GB diverse text) with full training transparency. Represents canonical standard MHA trained on general-purpose data.
Phi-2: Microsoft Research model, 2.7B parameters, 32 layers, 32 attention heads, 2560 hidden dimensions. Standard MHA trained on curated textbook-quality and code data. Represents optimal training conditions for geometric clarity.
Llama-3.2-1B: Meta model, 16 layers, 32 query heads, 8 key-value heads (4:1 grouped-query attention), 2048 hidden dimensions with rotary position embeddings. Trained on large-scale web data. Represents efficiency-optimized architecture for production deployment.
@@ 4.2. Geometric Extraction Protocol
We extract value manifolds, key orthogonality structure, and attention—-entropy trajectories using a uniform protocol across all models. All forward passes use each model’s native tokenizer and positional—-embedding scheme without modification.
@@
To avoid selection bias, we adopt a reproducible stratified—-entropy sampling procedure. We first generate 1,000 candidate prompts from five heterogeneous corpora (Wikipedia articles, news, fiction excerpts, code repositories, and general—-knowledge QA). For each prompt we compute the model’s next—-token entropy on the final position and partition the candidates into quintiles. We then uniformly sample $15$ prompts per quintile (total $75$ mixed—-domain prompts). For domain—-restricted experiments (e.g., mathematics, coding, philosophy), we filter the candidate pool to the relevant domain and apply the same stratified procedure. This ensures that geometric results do not depend on hand—-chosen or manually curated prompts.
@@
    the next—-token probability distribution $p(x_{T+1}\mid x_{1:T})$.
@@
For each model, we adopt a canonical PCA protocol on the final-layer value space. For a prompt of length $T$, we extract the value vectors $v^{(L)}_{T,h}\in\mathbb{R}^{d_{v}}$ for the final input token $T$ at the last layer $L$ from all attention heads $h=1,\dots,H$. We then concatenate the head-wise values into a single vector $\tilde{v}^{(L)}_{T}\in\mathbb{R}^{Hd_{v}}$ per prompt, so that each prompt contributes one $Hd_{v}$-dimensional sample. Before PCA, we standardize each coordinate across the prompt batch to zero mean and unit variance. All reported $\mathrm{PC}_{1}$ and $\mathrm{PC}_{1}{+}\mathrm{PC}_{2}$ statistics are computed from this standardized covariance matrix, separately for mixed-domain and domain-restricted prompt sets. Unless otherwise noted, PCA is run independently per model and per layer; cross-model analyses in Section Section 5.8 use a global PCA basis constructed by concatenating standardized value vectors across models.
@@
For each layer $\ell$ and head $h$, we take the key projection matrix $W^{(\ell,h)}_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$, extract its $d_{k}$ column vectors, and $\ell_{2}$-normalize each column to obtain unit vectors $\{\hat{k}_{i}\}_{i=1}^{d_{k}}\subset\mathbb{R}^{d_{\text{model}}}$. We then compute the mean off-diagonal absolute cosine similarity
@@ 4.3. In-Context Bayesian Updating Task
To test whether production models perform Bayesian updating during inference, we designed a controlled in-context learning task. Each prompt contains $k$ labeled sentiment examples (e.g., “happy is positive”, “sad is negative”) followed by a query word. We compute analytical Bayesian posteriors using a simple generative model with likelihood ratio 0.9:0.1 for consistent vs. inconsistent labels, generating 250 prompts across $k\in\{0,1,2,4,8\}$ with varying label imbalances.
@@ 4.4. Validation Criteria
Models meeting all three criteria exhibit Bayesian geometric signatures. Partial validation (2/3 criteria) suggests preserved qualitative structure with reduced clarity.
Threshold justification. These thresholds are anchored to quantitative baselines from controlled wind-tunnel settings and random-initialization controls. In bijection and HMM tasks, trained models produce nearly one-dimensional value manifolds with PC1 = 84–90% of variance explained. Under mixed-domain prompting in production models, multiple inference modes are simultaneously active, so we conservatively treat PC1+PC2 values in the 20–40% range as non-trivial structure over the random baseline of $\sim$5% for PC1. When prompts are domain-restricted, all models recover the same collapsed geometry as in wind-tunnel tasks (PC1 $\approx$ 0.75–0.85; PC1+PC2 $\approx$ 0.85–0.95).
For key orthogonality, random Gaussian $W_{K}$ matrices yield mean off-diagonal cosine values around 0.40–0.45. Trained models consistently achieve 0.03–0.20 depending on layer depth, so we adopt << 0.20 as a minimal criterion for structured hypothesis frames.
Wind-tunnel attention mechanisms reduce entropy by 85–90%. Production models face more heterogeneous workloads and architectural constraints (e.g., GQA), therefore we use a lenient 30% threshold to require meaningful focusing without expecting full collapse.
These thresholds are conservative: varying them within reasonable ranges does not change any model’s qualitative classification or the cross-model trends reported in this paper.
@@ 4.5. Statistical Validation
    Initialization baseline. For models where initialization checkpoints are publicly available (e.g., Pythia), we measure the mean off-diagonal cosine similarity of $W_{K}$ at step 0. These values (typically 0.350.35—-$0.45$, depending on the initialization scheme) reflect correlations introduced by weight initialization rather than pure Gaussian randomness.
@@
For each model, we compute per-head, per-position attentional entropies and average across heads and positions. Layerwise entropy reduction is evaluated relative to the entropy of the bottommost attention layer. All reductions exceeding 30% are significant under paired comparisons across prompts ($p<0.01$), and the architecture-dependent patterns (Section Section 5) are robust under bootstrap resampling.

### v3

@@ Abstract.
Paper I establishes that neural sequence models can implement exact Bayesian inference when they realize the required inference primitives: belief accumulation, belief transport, and random-access binding. Paper II shows that gradient descent learns these primitives through dynamics exhibiting a mechanistic analogy to EM, sculpting a characteristic geometry—low-dimensional value manifolds and progressively orthogonal keys. We investigate whether this geometric substrate persists in production-grade language models. Across Pythia, Phi-2, Llama-3, and Mistral families, we find that last-layer value representations organize along a single dominant axis whose position strongly correlates with predictive entropy, and that domain-restricted prompts collapse this structure into the same low-dimensional manifolds observed in wind-tunnel settings. This suggests that the geometric mechanisms underlying the inference primitives are preserved at scale. To probe the role of this geometry, we perform targeted interventions on the entropy-aligned axis of Pythia-410M during in-context learning. Removing or perturbing this axis selectively disrupts the local uncertainty geometry, whereas matched random-axis interventions leave it intact. However, these single-layer manipulations do not produce proportionally specific degradation in Bayesian-like behavior, indicating that the geometry is a privileged readout of uncertainty rather than a singular computational bottleneck. Taken together, our results show that modern language models preserve the geometric substrate that underlies the inference primitives, and organize their approximate Bayesian updates along this substrate.
@@ 1. Introduction
Paper I of this trilogy decomposes Bayesian inference into three inference primitives: belief accumulation (integrating evidence into a running posterior), belief transport (propagating beliefs through stochastic dynamics), and random-access binding (retrieving hypotheses by content). Different architectures realize different subsets: Transformers realize all three; Mamba realizes accumulation and transport; LSTMs realize only accumulation of static sufficient statistics; MLPs realize none. This taxonomy explains empirical performance across tasks. Paper II shows that gradient descent learns to implement these primitives through dynamics exhibiting a mechanistic analogy to EM, sculpting a characteristic geometry: value manifolds ordered by entropy, orthogonal key frames defining hypothesis directions, and layerwise attention sharpening implementing a geometric Bayes rule. These findings established that transformers can implement Bayesian inference when trained on tasks with known posteriors, and identified the geometric substrate that supports the inference primitives. What remains open is whether the same geometric mechanisms persist in large, naturally trained LLMs, where ground-truth posteriors are unavailable.
@@
Throughout this trilogy, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—e.g., filtering posteriors over hidden states—not a posterior over network weights. This is a statement about the function the transformer computes, not about weight-space uncertainty. Several factors make this question non-trivial:
@@
  • production models employ architectural optimizations (GQA, RoPE, sliding-window attention) absent from wind-tunnel setups;
@@
Rather than attempting to define Bayesian posteriors for natural language, we test whether the geometric substrate identified in Papers I–II persists across architectures and training regimes. We treat these geometric signatures as invariants: if transformers rely on similar computational principles at scale, the same value–key–attention geometry should appear even when exact posteriors cannot be measured.
@@
First, domain restriction produces a decisive bridge. Under mixed-domain prompts, value manifold dimensionality varies substantially across architectures (PC1+PC2 ranging from $\sim$15% in Mistral to $\sim$99% in Pythia-410M), reflecting different inductive biases and training regimes. However, single-domain prompts consistently collapse the manifold toward one dimension (PC1+PC2 $\approx$ 70–95%), approaching the geometric regime observed in wind-tunnel experiments. This collapse shows that production LLMs contain the same entropy-ordered Bayesian axis that wind-tunnel transformers learn explicitly. Second, Bayesian updating persists at inference time. In a controlled in-context learning experiment (SULA), models move smoothly along their value manifold as more evidence is supplied, and manifold coordinates correlate strongly with analytical Bayesian entropy. This demonstrates that the geometry is not merely a training artifact - it is used during inference. Third, static and dynamic geometric signatures separate cleanly. Value manifolds and key orthogonality are universal across architectures, including sliding-window and MoE variants. Dynamic attention focusing, however, depends on routing capacity: strong in full-sequence MHA, moderate in GQA, and weak or noisy in Mistral. This matches the frame–precision dissociation predicted in Paper II.
@@
Paper I established which architectures can implement Bayesian inference, taxonomized by three inference primitives (belief accumulation, belief transport, random-access binding). Paper II showed how gradient dynamics learn to implement these primitives through mechanisms exhibiting a mechanistic analogy to EM, sculpting characteristic geometric structures. However, these results were obtained in synthetic domains with analytically specified likelihoods. The present work answers a distinct question: do production-grade LLMs, trained on heterogeneous natural language, spontaneously develop the same geometric signatures that underlie the inference primitives? We show that the three hallmarks of Bayesian geometry—low-dimensional value manifolds, orthogonal hypothesis frames, and evidence-dependent movement along entropy-aligned directions—persist across four model families (Pythia, Phi-2, Llama-3, Mistral). Moreover, we provide the first large-scale evidence that these structures are functionally engaged during inference in a naturalistic task (SULA), even when their causal role is distributed rather than bottlenecked. This establishes that the geometric substrate for the inference primitives is a stable inductive bias of modern transformers, not an artifact of synthetic tasks.
@@
    Domain-restriction bridge. When prompts are restricted to a coherent domain, the value manifold collapses to one or two principal components explaining 80–95% of variance, numerically matching the geometric regime predicted by Paper I and derived in Paper II.
@@ 2. Background: Bayesian Geometry in Controlled Settings
Papers I–II established that small transformers trained in controlled “Bayesian wind-tunnel” settings perform near-exact Bayesian inference on tasks with analytically tractable posteriors. We summarize only the key findings needed for the production-model analysis; full experimental details appear in those papers.
@@
Two families of synthetic tasks provided ground-truth posteriors: (1) Bijection learning (Paper I): models infer a random bijection $\pi:V\to W$ from in-context examples. Because the hypothesis space has size $K!$ (e.g., $3.6\times 10^{6}$ for $K=10$), memorization is impossible; exact analytic posterior trajectories are available. Transformers achieve $\text{MAE}<0.1$ bits between model and Bayes-optimal predictive entropy. (2) HMM filtering (Paper I): models track posterior distributions over latent states in Hidden Markov Models with $S=5$ states and $V=5$ emissions. Transformers match analytic posteriors with KL divergence <0.05<0.05 bits, including strong length generalization.
@@
Paper II showed that cross-entropy gradients generate this geometry via coupled specialization of queries, keys, and values. A predicted frame–precision dissociation emerges: attention patterns (the frame) stabilize early, while value manifolds (precision) continue refining.
@@
In this paper, we evaluate whether the static and dynamic geometric signatures identified in wind-tunnel models - entropy-ordered value manifolds, orthogonal hypothesis frames, and layerwise attention refinement - persist in production-scale LLMs and whether these structures are used during inference. Section 5 presents our empirical findings across architectures.
@@ 4.1. Model Selection
We selected three production models to span architectural and training variations: Pythia-410M (Biderman et al., 2023): GPT-NeoX architecture, 24 layers, 16 attention heads, 1024 hidden dimensions. Trained on the Pile corpus (800GB diverse text) with full training transparency. Represents canonical standard MHA trained on general-purpose data. Phi-2: Microsoft Research model, 2.7B parameters, 32 layers, 32 attention heads, 2560 hidden dimensions. Standard MHA trained on curated textbook-quality and code data. Represents optimal training conditions for geometric clarity. Llama-3.2-1B: Meta model, 16 layers, 32 query heads, 8 key-value heads (4:1 grouped-query attention), 2048 hidden dimensions with rotary position embeddings. Trained on large-scale web data. Represents efficiency-optimized architecture for production deployment.
@@ 4.2. Geometric Extraction Protocol
We extract value manifolds, key orthogonality structure, and attention–entropy trajectories using a uniform protocol across all models. All forward passes use each model’s native tokenizer and positional–embedding scheme without modification.
@@
To avoid selection bias, we adopt a reproducible stratified–entropy sampling procedure. We first generate 1,000 candidate prompts from five heterogeneous corpora (Wikipedia articles, news, fiction excerpts, code repositories, and general–knowledge QA). For each prompt we compute the model’s next–token entropy on the final position and partition the candidates into quintiles. We then uniformly sample $15$ prompts per quintile (total $75$ mixed–domain prompts). For domain–restricted experiments (e.g., mathematics, coding, philosophy), we filter the candidate pool to the relevant domain and apply the same stratified procedure. This ensures that geometric results do not depend on hand–chosen or manually curated prompts.
@@
    the next–token probability distribution $p(x_{T+1}\mid x_{1:T})$.
@@
For each model, we adopt a canonical PCA protocol on the final-layer value space. For a prompt of length $T$, we extract the V-projection outputs $v^{(L)}_{T,h}=W_{V}^{(h)}x_{T}\in\mathbb{R}^{d_{v}}$ for the final input token $T$ at the last layer $L$ from all attention heads $h=1,\dots,H$. We then concatenate the head-wise values into a single vector $\tilde{v}^{(L)}_{T}\in\mathbb{R}^{Hd_{v}}$ per prompt, so that each prompt contributes one $Hd_{v}$-dimensional sample. Before PCA, we standardize each coordinate across the prompt batch to zero mean and unit variance. All reported $\mathrm{PC}_{1}$ and $\mathrm{PC}_{1}{+}\mathrm{PC}_{2}$ statistics are computed from this standardized covariance matrix, separately for mixed-domain and domain-restricted prompt sets. Unless otherwise noted, PCA is run independently per model and per layer; cross-model analyses in Section 5.8 use a global PCA basis constructed by concatenating standardized value vectors across models.
@@
For each layer $\ell$ and head $h$, we take the key projection matrix $W^{(\ell,h)}_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$, extract its $d_{k}$ column vectors, and $\ell_{2}$-normalize each column to obtain unit vectors $\{\hat{k}_{i}\}_{i=1}^{d_{k}}\subset\mathbb{R}^{d_{\text{model}}}$. Note that we measure orthogonality of the learned parameter matrix columns, not token-conditioned key vectors; this captures the static geometric structure of the hypothesis frames established during training. We then compute the mean off-diagonal absolute cosine similarity
@@ 4.3. In-Context Bayesian Updating Task
To test whether production models perform Bayesian updating during inference, we designed a controlled in-context learning task. Each prompt contains $k$ labeled sentiment examples (e.g., “happy is positive”, “sad is negative”) followed by a query word. We compute analytical Bayesian posteriors using a Beta-Bernoulli model with uniform Beta(1,1) prior and likelihood ratio 0.9:0.1 for consistent vs. inconsistent labels, generating 250 prompts across $k\in\{0,1,2,4,8\}$ with varying label imbalances. SULA is not intended as a model of natural-language reasoning, but as a controlled probe isolating Bayesian updating from linguistic ambiguity.
@@ 4.4. Validation Criteria
Models meeting all three criteria exhibit Bayesian geometric signatures. Partial validation (2/3 criteria) suggests preserved qualitative structure with reduced clarity. Threshold justification. These thresholds are anchored to quantitative baselines from controlled wind-tunnel settings and random-initialization controls. In bijection and HMM tasks, trained models produce nearly one-dimensional value manifolds with PC1 = 84–90% of variance explained. Under mixed-domain prompting in production models, multiple inference modes are simultaneously active, so we conservatively treat PC1+PC2 values in the 20–40% range as non-trivial structure over the random baseline of $\sim$5% for PC1. When prompts are domain-restricted, all models recover the same collapsed geometry as in wind-tunnel tasks (PC1 $\approx$ 0.75–0.85; PC1+PC2 $\approx$ 0.85–0.95). For key orthogonality, random Gaussian $W_{K}$ matrices yield mean off-diagonal cosine values around 0.40–0.45. Trained models consistently achieve 0.03–0.20 depending on layer depth, so we adopt << 0.20 as a minimal criterion for structured hypothesis frames. Wind-tunnel attention mechanisms reduce entropy by 85–90%. Production models face more heterogeneous workloads and architectural constraints (e.g., GQA), therefore we use a lenient 30% threshold to require meaningful focusing without expecting full collapse. These thresholds are conservative: varying them within reasonable ranges does not change any model’s qualitative classification or the cross-model trends reported in this paper.
@@ 4.5. Statistical Validation
    Initialization baseline. For models where initialization checkpoints are publicly available (e.g., Pythia), we measure the mean off-diagonal cosine similarity of $W_{K}$ at step 0. These values (typically 0.350.350.450.45, depending on the initialization scheme) reflect correlations introduced by weight initialization rather than pure Gaussian randomness.
@@
For each model, we compute per-head, per-position attentional entropies and average across heads and positions. Layerwise entropy reduction is evaluated relative to the entropy of the bottommost attention layer. All reductions exceeding 30% are significant under paired comparisons across prompts ($p<0.01$), and the architecture-dependent patterns (Section 5) are robust under bootstrap resampling.